OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? 600 60 300 120 <-- this is what i got I'm not sure though please correct if I'm wrong

OpenStudy (dmb1215):

i will second that

OpenStudy (anonymous):

thank you @V1r3z

OpenStudy (anonymous):

My bad it was 600.

OpenStudy (anonymous):

A) 600.

OpenStudy (anonymous):

06x+24=.05x+30 .06x=.05x+6 subtract 24 both sides .01x=6 subtract .05 both sides :) x=6/.01 which is 600 when you divide.

OpenStudy (anonymous):

oh okay thank you @V1r3z